# The Uniform Mordell–Lang Conjecture

### v3

@@ 5.1. A result in family
  • and each irreducible subvariety $X$ of $A$ defined over $\overline{\mathbb{Q}}$ with $\dim X=r$ and $\deg_{L}X=d$, such that $X$ generates $A$,
@@
(6.1) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(P)\leq c_{1}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}$
where $L_{-}:=[-1]^{*}L$, is contained in a proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}X^{\prime}<c_{2}$.
@@
Let $(A,L)$ and $X$ be as in the proposition.
By Proposition 4.1, there exist constants $c_{1}^{\prime}=c_{1}^{\prime}(g,l,r,d)>0$ ,$c_{2}^{\prime}=c_{2}^{\prime}(g,l,r,d)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(g,l,r,d)>0$ such that
@@
We split up to two cases on whether $\max\{1,h_{\mathrm{Fal}}(A)\}\leq\max\{1,2c_{3}^{\prime}/c_{1}^{\prime}\}$.
@@
$\Sigma:=\left\{P\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(P)\leq c_{1}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}$
@@ 7.1. A theorem of Rémond
(7.2) #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X,\deg_{L}A)^{\mathrm{rk}\Gamma+1}.$
@@
We start with a dévissage. More precisely, we reduce to the case where
(Hyp): $X$ generates $A$.
Indeed, let $A^{\prime}$ be the abelian subvariety of $A$ generated by $X-X$. Then $X\subseteq A^{\prime}+Q$ for some $Q\in A(\overline{\mathbb{Q}})$. The subgroup $\Gamma^{\prime}$ of $A(\overline{\mathbb{Q}})$ generated by $\Gamma$ and $Q$ has rank $\leq\mathrm{rk}\Gamma+1$. We have $(X-Q)^{\circ}=X^{\circ}-Q$ by definition of the Ueno locus, $(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma\subseteq(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma^{\prime}=X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\prime}$ and $\deg_{L}(X-Q)=\deg_{L}X$. By Lemma 2.5, (Hyp) yields $\deg_{L}A^{\prime}\leq c^{\prime}(g,\deg_{L}X)$. Therefore if (7.2) holds true for $X-Q\subseteq A^{\prime}$, $L|_{A^{\prime}}$ and $\Gamma^{\prime}\cap A^{\prime}(\overline{\mathbb{Q}})$, then #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma^{\prime}+1}\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma+2}$. So we can conclude by replacing $c$ with $c^{2}$. Thus we are reduced to the case where (Hyp) holds true.
We take the formulation of David–Philippon [DP07, Thm.6.8] of Rémond’s result [Rém00a].
Let $c_{\mathrm{NT}}$ and $h_{1}$ be as from [DP07, Thm.6.8]. It is known that $c_{\mathrm{NT}},h_{1}\leq c^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}$ for some $c^{\prime}=c^{\prime}(g,\deg_{L}A)>0$; see [DP07, equation (6.41)].
@@
(7.3) #$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta\max\{1,c_{\mathrm{NT}},h_{1}\}\right\}\leq\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta c^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}.$
(Hyp) allows us to apply (7.1) to $X-P_{0}$ for each $P_{0}\in X(\overline{\mathbb{Q}})$.
@@
Consider the real vector space $\Gamma\otimes_{\mathbb{Q}}\mathbb{R}$ endowed with the Euclidean norm $|\cdot|=\hat{h}_{L}^{1/2}$. By an elementary ball packing argument, any subset of $\Gamma\otimes\mathbb{R}$ contained in a closed ball of radius $R$ centered at 00 is covered by at most $(1+2R/r)^{\mathrm{rk}\Gamma}$ closed balls of radius $r$ centered at the elements $P-P_{0}$ of the given subset (7.1); see [Rém00a, Lem.6.1]. Thus the number of balls in the covering is at most $(1+2\sqrt{\eta c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}$. But each closed ball of radius $r$ centered at some $P-P_{0}$ in (7.1) contains at most $c$ elements by (7.1). So
(7.4) #$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta c^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{0}(1+2\sqrt{\eta c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}.$
@@
(7.5) #$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta\max\{1,c_{\mathrm{NT}},h_{1}\}\right\}\leq c_{0}(1+2\sqrt{\eta c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}.$
@@
(7.6) #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq(c^{\prime\prime})^{\mathrm{rk}\Gamma+1}\cdot c_{0}(1+2\sqrt{c^{\prime\prime}c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}$
@@
Let $A$ be an abelian variety of dimension $g$ and $L$ be an ample line bundle on $A$. Then with $L$ replace by $L\otimes L_{-}$, (Hyp pack) holds true for each irreducible subvariety $X$ of $A$ which generates $A$ and each finite rank subgroup $\Gamma$ of $A(\overline{\mathbb{Q}})$.
@@
For an arbitrary $r\geq 1$. Assume the theorem is proved for $10,1,\ldots,r-$ 1.
@@
Apply Theorem 1.2 to $X-P_{0}$. Then we have constants $c_{1}=c_{1}(g,d)>0$ and $c_{2}=c_{2}(g,d)>0$ such that for $\left\{P-P_{0}\in X^{\circ}(\overline{\mathbb{Q}})-P_{0}:\hat{h}_{L\otimes L_{-}}(P-P_{0})\leq c_{1}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}$ is contained in a proper Zariski closed $X^{\prime}\subsetneq X-P_{0}$ with $\deg_{L}X^{\prime}<c_{2}$. In particular, the number of irreducible components of $X^{\prime}$ is $<c_{2}$.
@@
(7.7) #$\#(X^{\dagger}-Q)^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\dagger}\leq(c^{\dagger})^{\mathrm{rk}\Gamma^{\dagger}+1}\leq(c^{\dagger})^{\mathrm{rk}\Gamma+2}$
@@
(7.8) #$\#(X^{\dagger})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\dagger}\leq(c^{\dagger})^{\mathrm{rk}\Gamma+2}$
@@
(7.9) #$\#(X^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c_{2}\max_{X^{\dagger}}\{c^{\dagger}\}^{\mathrm{rk}\Gamma+2}\leq c_{3}^{\mathrm{rk}\Gamma+1}$
@@
Notice that $\deg_{L\otimes L_{-}}X=2^{\dim X}\deg_{L}X\leq 2^{g}d$ and $\deg_{L\otimes L_{-}}A=2^{g}\deg_{L}A=2^{g}l$.
@@
As before, we start by reducing to the case where $X$ generates $A$. Indeed, let $A^{\prime}$ be the abelian subvariety of $A$ generated by $X-X$. Then $X\subseteq A^{\prime}+Q$ for some $Q\in A(\overline{\mathbb{Q}})$. The subgroup $\Gamma^{\prime}$ of $A(\overline{\mathbb{Q}})$ generated by $\Gamma$ and $Q$ has rank $\leq\mathrm{rk}\Gamma+1$. We have $(X-Q)^{\circ}=X^{\circ}-Q$ by definition of the Ueno locus, $(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma\subseteq(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma^{\prime}=X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\prime}$ and $\deg_{L}(X-Q)=\deg_{L}X$. If (1.2) holds true for $X-Q\subseteq A^{\prime}$, $L|_{A^{\prime}}$ and $\Gamma^{\prime}\cap A^{\prime}(\overline{\mathbb{Q}})$, then #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,d)^{\mathrm{rk}\Gamma^{\prime}+1}\leq c(g,d)^{\mathrm{rk}\Gamma+2}$. So we can conclude by replacing $c$ with $c^{2}$. Thus we are reduced to the case where $X$ generates $A$. In particular, we have $l\leq c^{\prime}(g,d)$ by Lemma 2.5.
By Proposition 7.2 and Theorem 7.1, we have #$\#X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,l,d)^{\mathrm{rk}\Gamma+1}$. Thus (1.2) holds true because $l\leq c^{\prime}(g,d)$. Hence we are done. ∎
7.3. From $\overline{\mathbb{Q}}$ to arbitrary $F$ of characteristic 00
@@
Assume Theorem 1.1 holds true for $F=\overline{\mathbb{Q}}$. Then Theorem 1.1 holds true, under the extra assumptions that $\Gamma$ is finitely generated, for arbitrary $F$ of characteristic 00 with $F=\overline{F}$.
@@
Now let $F$ be an arbitrary algebraic closed field of characteristic 00. Let $A$, $L$, $X$ and $\Gamma$ be as in Theorem 1.1 with $\Gamma$ finitely generated. Write $\rho=\mathrm{rk}\Gamma$. Let $\gamma_{1},\ldots,\gamma_{r}\in A(F)$ be generators of $\Gamma$ with $\gamma_{\rho+1},\ldots,\gamma_{r}$ torsion.
@@
  • $A$ extends to an abelian scheme $\mathcal{A}\rightarrow V$ of relative dimension $g$,
@@
Moreover, up to replacing $V$ by a Zariski open dense subset, we may choose $\mathfrak{X}$ over $V$ to be a “Good model” as explained in [Maz00, pp.219–220] (which uses [Hin88, App.1, Lemma A]) such that the Ueno locus of $\mathfrak{X}_{v}$ is the specialization of the Ueno locus of $X$ for each $v\in V(\overline{\mathbb{Q}})$. Hence $\mathfrak{X}_{v}^{\circ}$ is the specialization of $X^{\circ}$.
@@
Let $v\in\Theta$. Then #$\#X^{\circ}(F)\cap\Gamma\leq\#\mathfrak{X}_{v}^{\circ}(\overline{\mathbb{Q}})\cap\Gamma_{v}$. By the result over $\overline{\mathbb{Q}}$, we have #$\#\mathfrak{X}_{v}^{\circ}(\overline{\mathbb{Q}})\cap\Gamma_{v}\leq c(g,\deg_{{\mathcal{L}}_{v}}\mathfrak{X}_{v})^{1+\mathrm{rk}\Gamma_{v}}\leq c(g,\deg_{L}X)^{1+\rho}$. Hence we are done. ∎
@@
Let $F$ be an arbitrary algebraic closed field of characteristic 00. Let $A$, $L$ and $X$ be as in Theorem 1.1. Let $\Gamma$ be a subgroup of $A(F)$ of finite rank $\rho$.
@@
(7.11) #$\#X^{\circ}(F)\cap\frac{1}{n}\Gamma_{0}\leq c^{1+\rho}.$
@@
The union above canbe expressed in a quantitative way. First, by Bogomolov [Bog81, Thm. 1], each $B\in\Sigma(X;A)$ satisfies $\deg_{L}B\leq c_{3}$ for some constant $c_{3}=c_{3}(\dim A,\deg_{L}X)>0$, and hence #$\#\Sigma(X;A)\leq c_{4}=c_{4}(\dim A,\deg_{L}X)$ by [Rém00a, Prop. 4.1]; here we used Lemma 2.5. Next, $X_{B}$ can be constructed as follows. Let $B^{\perp}$ be a complement of $B$, i.e. $B\cap B^{\perp}$ is finite and $B+B^{\perp}=A$. It is possible to choose such a $B^{\perp}$ with $\deg_{L}B^{\perp}\leq c_{5}^{\prime}(g,\deg_{L}A,\deg_{L}B)$; see [MW93]. Then we can choose $X_{B}:=\bigcap_{b\in B(F)}(X-b)\bigcap B^{\perp}$. Notice that by dimension reasons, this intersection must be a finite intersection of at most $\dim X\leq\dim A$ members. So $\deg_{L}X_{B}\leq c_{5}(\dim A,\deg_{L}X)$ by Bézout’s Theorem and Lemma 2.5. In particular $X_{B}$ has $\leq c_{5}$ irreducible components $X_{B,1},\ldots,X_{B,m_{B}}$.
@@
Let us bound $n_{B}$ for each $B\in\Sigma(X;A)$. Applying Theorem 1.1 to each irreducible component $X_{B,k}$ of $X_{B}$, we get #$\#X_{B,k}^{\circ}(F)\cap\Gamma\leq c^{1+\mathrm{rk}\Gamma}$ for some $c=c(\dim A,\deg_{L}X_{B,k})>0$. But we have seen that $X_{B}$ has $\leq c_{5}$ components and that $\deg_{L}X_{B,k}\leq\deg_{L}X_{B}\leq c_{5}$. So #$\#X_{B}^{\circ}(F)\cap\Gamma$, and hence $n_{B}$, is $\leq c_{6}(\dim A,\deg_{L}X)^{1+\mathrm{rk}\Gamma}$.
@@
For arbitrary $r$, assume the theorem is proved for $10,\ldots,r-$ 1.
Let $A^{\prime}$ be the abelian subvariety of $A$ generated by $X-X$, then $\deg_{L}A^{\prime}\leq cd^{2g}$ by Lemma 2.5. We can apply Proposition 5.1 to $X$ and $(A^{\prime},L|_{A^{\prime}})$ to conclude that the set

### v4

@@ 5.1. A preliminary result for the restricted Hilbert scheme
  • and each irreducible subvariety $X$ of $A$ with $\dim X=r$ and $\deg_{L}X=d$, such that $X$ generates $A$,
@@
(6.1) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes[-1]^{*}L}(P)\leq c_{1}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}$
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}X^{\prime}<c_{2}$.
@@
Let $(A,L)$ and $X$ be as in the proposition. Denote for simplicity by $L_{-}:=[-1]^{*}L$.
By Proposition 4.1, there exist constants $c_{1}^{\prime}=c_{1}^{\prime}(g,l,r,d)>0$ $c_{2}^{\prime}=c_{2}^{\prime}(g,l,r,d)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(g,l,r,d)>0$ such that
@@
We split up to two cases on whether or not $\max\{1,h_{\mathrm{Fal}}(A)\}\leq\max\{1,2c_{3}^{\prime}/c_{1}^{\prime}\}$.
@@
$\Sigma:=\left\{P\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes[-1]^{*}L}(P)\leq c_{1}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}$
@@ 7.1. A theorem of Rémond
(7.2) $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X,\deg_{L}A)^{\mathrm{rk}\Gamma+1}.$
@@
We start with a dévissage. More precisely, we reduce to the case where $X$ generates $A$. Indeed, let $A^{\prime}$ be the abelian subvariety of $A$ generated by XXX-X. Then $X\subseteq A^{\prime}+Q$ for some $Q\in A(\overline{\mathbb{Q}})$. The subgroup $\Gamma^{\prime}$ of $A(\overline{\mathbb{Q}})$ generated by $\Gamma$ and $Q$ has rank $\leq\mathrm{rk}\Gamma+1$. We have $(X-Q)^{\circ}=X^{\circ}-Q$ by definition of the Ueno locus, $(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma\subseteq(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma^{\prime}=X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\prime}$ and $\deg_{L}(X-Q)=\deg_{L}X$. By Lemma 2.5, we have $\deg_{L}A^{\prime}\leq c^{\prime}(g,\deg_{L}X)$. Therefore if (7.2) holds true for $X-Q\subseteq A^{\prime}$, $L|_{A^{\prime}}$ and $\Gamma^{\prime}\cap A^{\prime}(\overline{\mathbb{Q}})$, then $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma^{\prime}+1}\leq c(g,\deg_{L}X)^{\mathrm{rk}\Gamma+2}$. So we can conclude by replacing $c$ with $c^{2}$. Thus we are reduced to the case where $X$ generates $A$.
From now on, assume $X$ generates $A$. We take the formulation of David–Philippon [DP07, Thm. 6.8] of Rémond’s result [Rém00a].
It follows from Tate’s construction that there exists a constant $c_{\mathrm{NT}}(A)\geq 0$, which depends on $A$, such that $|\hat{h}_{L}(P)-h(P)|\leq c_{\mathrm{NT}}(A)$ for all $P\in A(\overline{\mathbb{Q}})$.
Let $h_{1}(A)$ denote the Weil height of the polynomials defining the addition and the substraction on $A$.
It is known that $c_{\mathrm{NT}}(A),h_{1}(A)\leq c^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}$ for some $c^{\prime}=c^{\prime}(g,\deg_{L}A)>0$; see [DP07, equation (6.41)]. Alternatively this can be deduced from [DGH21, (8.4) and (8.7)].
@@
(7.3) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta\max\{1,c_{\mathrm{NT}}(A),h_{1}(A)\}\right\}\\ \leq\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta c^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}.$
Since $X$ generates $A$, we can apply (7.1) to $X-P_{0}$ for each $P_{0}\in X(\overline{\mathbb{Q}})$.
@@
Consider the real vector space $\Gamma\otimes_{\mathbb{Q}}\mathbb{R}$ endowed with the Euclidean norm $|\cdot|=\hat{h}_{L}^{1/2}$. By an elementary ball packing argument, any subset of $\Gamma\otimes\mathbb{R}$ contained in a closed ball of radius $R$ centered at 0 is covered by at most $(1+2R/r)^{\mathrm{rk}\Gamma}$ closed balls of radius $r$ centered at the elements $P-P_{0}$ of the given subset (7.1); see [Rém00a, Lem. 6.1]. Thus the number of balls in the covering is at most $(1+2\sqrt{\eta c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}$. But each closed ball of radius $r$ centered at some $P-P_{0}$ in (7.1) contains at most $c$ elements by (7.1). So
(7.4) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta c^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}\leq c_{0}(1+2\sqrt{\eta c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}.$
@@
(7.5) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma:\hat{h}_{L}(P)\leq\eta\max\{1,c_{\mathrm{NT}}(A),h_{1}(A)\}\right\}\leq c_{0}(1+2\sqrt{\eta c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}.$
@@
(7.6) $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq(c^{\prime\prime})^{\mathrm{rk}\Gamma+1}\cdot c_{0}(1+2\sqrt{c^{\prime\prime}c^{\prime}c_{0}})^{\mathrm{rk}\Gamma}$
@@
Let $A$ be an abelian variety of dimension $g$ and $L$ be an ample line bundle on $A$. Then with $L$ replaced by $L\otimes[-1]^{*}L$, (Hyp pack) holds true for each irreducible subvariety $X$ of $A$ which generates $A$ and each finite rank subgroup $\Gamma$ of $A(\overline{\mathbb{Q}})$.
@@
For an arbitrary $r\geq 1$. Assume the theorem is proved for $0,1,\ldots,r-1$.
@@
Apply Theorem 1.2 to $X-P_{0}$. Then we have constants $c_{1}=c_{1}(g,d)>0$ and $c_{2}=c_{2}(g,d)>0$ such that for $\left\{P-P_{0}\in X^{\circ}(\overline{\mathbb{Q}})-P_{0}:\hat{h}_{L\otimes L_{-}}(P-P_{0})\leq c_{1}\max\{1,h_{\mathrm{Fal}}(A)\}\right\}$ is contained in a proper Zariski closed $X^{\prime}\subsetneq X-P_{0}$ with $\deg_{L}X^{\prime}<c_{2}$. In particular, the number of irreducible components of $X^{\prime}$ is $c_{2}$.
@@
(7.7) $(X^{\dagger}-Q)^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\dagger}\leq(c^{\dagger})^{\mathrm{rk}\Gamma^{\dagger}+1}\leq(c^{\dagger})^{\mathrm{rk}\Gamma+2}$
@@
(7.8) $(X^{\dagger})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\dagger}\leq(c^{\dagger})^{\mathrm{rk}\Gamma+2}$
@@
(7.9) $(X^{\prime})^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c_{2}\max_{X^{\dagger}}\{c^{\dagger}\}^{\mathrm{rk}\Gamma+2}\leq c_{3}^{\mathrm{rk}\Gamma+1}$
@@
Notice that $\deg_{L\otimes[-1]^{*}L}X=2^{\dim X}\deg_{L}X\leq 2^{g}d$ and $\deg_{L\otimes[-1]^{*}L}A=2^{g}\deg_{L}A=2^{g}l$.
@@
As before, we start by reducing to the case where $X$ generates $A$. Indeed, let $A^{\prime}$ be the abelian subvariety of $A$ generated by XXX-X. Then $X\subseteq A^{\prime}+Q$ for some $Q\in A(\overline{\mathbb{Q}})$. The subgroup $\Gamma^{\prime}$ of $A(\overline{\mathbb{Q}})$ generated by $\Gamma$ and $Q$ has rank $\leq\mathrm{rk}\Gamma+1$. We have $(X-Q)^{\circ}=X^{\circ}-Q$ by definition of the Ueno locus, $(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma\subseteq(X^{\circ}(\overline{\mathbb{Q}})-Q)\cap\Gamma^{\prime}=X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma^{\prime}$ and $\deg_{L}(X-Q)=\deg_{L}X$. If (1.2) holds true for $X-Q\subseteq A^{\prime}$, $L|_{A^{\prime}}$ and $\Gamma^{\prime}\cap A^{\prime}(\overline{\mathbb{Q}})$, then $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,d)^{\mathrm{rk}\Gamma^{\prime}+1}\leq c(g,d)^{\mathrm{rk}\Gamma+2}$. So we can conclude by replacing $c$ with $c^{2}$. Thus we are reduced to the case where $X$ generates $A$. In particular, we have $l\leq c^{\prime}(g,d)$ by Lemma 2.5.
By Proposition 7.2 and Theorem 7.1, we have $X^{\circ}(\overline{\mathbb{Q}})\cap\Gamma\leq c(g,l,d)^{\mathrm{rk}\Gamma+1}$. Thus (1.2) holds true because $l\leq c^{\prime}(g,d)$. Hence we are done. ∎
7.3. From $\overline{\mathbb{Q}}$ to arbitrary $F$ of characteristic 0
@@
Assume Theorem 1.1 holds true for $F=\overline{\mathbb{Q}}$. Then Theorem 1.1 holds true, under the extra assumptions that $\Gamma$ is finitely generated, for arbitrary $F$ of characteristic 0 with $F=\overline{F}$.
@@
Now let $F$ be an arbitrary algebraic closed field of characteristic 0. Let $A$, $L$, $X$ and $\Gamma$ be as in Theorem 1.1 with $\Gamma$ finitely generated. Write $\rho=\mathrm{rk}\Gamma$. Let $\gamma_{1},\ldots,\gamma_{r}\in A(F)$ be generators of $\Gamma$ with $\gamma_{\rho+1},\ldots,\gamma_{r}$ torsion.
@@
  • AA extends to an abelian scheme $\mathcal{A}\rightarrow V$ of relative dimension $g$,
@@
Moreover, up to replacing $V$ by a Zariski open dense subset, we may choose $\mathfrak{X}$ over $V$ to be a “Good model” as explained in [Maz00, p. 219–220] (which uses [Hin88, App. 1, Lemma A]) such that the Ueno locus of $\mathfrak{X}_{v}$ is the specialization of the Ueno locus of $X$ for each $v\in V(\overline{\mathbb{Q}})$. Hence $\mathfrak{X}_{v}^{\circ}$ is the specialization of $X^{\circ}$.
@@
Let $v\in\Theta$. Then $X^{\circ}(F)\cap\Gamma\leq\#\mathfrak{X}_{v}^{\circ}(\overline{\mathbb{Q}})\cap\Gamma_{v}$. By the result over $\overline{\mathbb{Q}}$, we have $\mathfrak{X}_{v}^{\circ}(\overline{\mathbb{Q}})\cap\Gamma_{v}\leq c(g,\deg_{{\mathcal{L}}_{v}}\mathfrak{X}_{v})^{1+\mathrm{rk}\Gamma_{v}}\leq c(g,\deg_{L}X)^{1+\rho}$. Hence we are done. ∎
@@
Let $F$ be an arbitrary algebraic closed field of characteristic 0. Let $A$, $L$ and $X$ be as in Theorem 1.1. Let $\Gamma$ be a subgroup of $A(F)$ of finite rank $\rho$.
@@
(7.11) $X^{\circ}(F)\cap\frac{1}{n}\Gamma_{0}\leq c^{1+\rho}.$
@@
The union above can be expressed in a quantitative way. First, by Bogomolov [Bog81, Thm. 1], each $B\in\Sigma(X;A)$ satisfies $\deg_{L}B\leq c_{3}$ for some constant $c_{3}=c_{3}(\dim A,\deg_{L}X)>0$, and hence $\Sigma(X;A)\leq c_{4}=c_{4}(\dim A,\deg_{L}X)$ by [Rém00a, Prop. 4.1]; here we used Lemma 2.5. Next, $X_{B}$ can be constructed as follows. Let $B^{\perp}$ be a complement of $B$, i.e. $B\cap B^{\perp}$ is finite and $B+B^{\perp}=A$. It is possible to choose such a $B^{\perp}$ with $\deg_{L}B^{\perp}\leq c_{5}^{\prime}(g,\deg_{L}A,\deg_{L}B)$; see [MW93]. Then we can choose $X_{B}:=\bigcap_{b\in B(F)}(X-b)\bigcap B^{\perp}$. Notice that by dimension reasons, this intersection must be the component of a finite intersection of at most $\dim X\leq\dim A$ members. So $\deg_{L}X_{B}\leq c_{5}(\dim A,\deg_{L}X)$ by Bézout’s Theorem and Lemma 2.5. In particular $X_{B}$ has $\leq c_{5}$ irreducible components $X_{B,1},\ldots,X_{B,m_{B}}$.
@@
Let us bound $n_{B}$ for each $B\in\Sigma(X;A)$. Applying Theorem 1.1 to each irreducible component $X_{B,k}$ of $X_{B}$, we get $X_{B,k}^{\circ}(F)\cap\Gamma\leq c^{1+\mathrm{rk}\Gamma}$ for some $c=c(\dim A,\deg_{L}X_{B,k})>0$. But we have seen that $X_{B}$ has $\leq c_{5}$ components and that $\deg_{L}X_{B,k}\leq\deg_{L}X_{B}\leq c_{5}$. So $X_{B}^{\circ}(F)\cap\Gamma$, and hence $n_{B}$, is $\leq c_{6}(\dim A,\deg_{L}X)^{1+\mathrm{rk}\Gamma}$.
@@
For arbitrary $r$, assume the theorem is proved for $0,\ldots,r-1$.
Let $A^{\prime}$ be the abelian subvariety of $A$ generated by XXX-X, then $\deg_{L}A^{\prime}\leq cd^{2g}$ by Lemma 2.5. We can apply Proposition 5.1 to $X$ and $(A^{\prime},L|_{A^{\prime}})$ to conclude that the set